Sketch the graph of the equation.

y =

5
x2 + 1


Identify the intercept.

(x, y) = (  )
Test for symmetry. (Select all that apply.)
Answer #1

Sketchy= bx²+1 y = 5uhti 2 x = Žly-1) - y = 5x31 upweord parabola with vertugas (0,1) 0,1 YU Intercept X-infcreept - No inter